### v3

Abstract
This paper proposes a novel approach for Fourier extension by reformulating the problem as solving a compact operator equation with perturbed data. To mitigate oscillations in the extended region, we introduce a weighted generalized inverse framework that shifts the approximation target from the best approximate solution to a weighted variant. Theoretical error bounds for the solution are established. We identify computational challenges arising from machine precision limitations when implementing the original weighted operator and propose a corrected operator to address this issue. The relevant parameters involved in the method are further tested, and finally the effectiveness of the method is verified through numerical experiments.
Keywords:
Fourier ExtensionsFourier ContinuationWeighted Generalized Inverse
MSC:
42A10 41A17 65T40 42C15
1 Introduction
Due to its flexibility, frame approximation has received increasing attention in recent years. Its related theories and numerical algorithms are being widely studied and have made substantial progress adcock2019frames ; adcock2020frames ; adcock2020approximating ; adcock2021frame ; herremans2024efficient ; ebner2023regularization . Fourier extension is an important form of frame approximation, which has significantly contributed to the advancement of the field of frame theory. Initially proposed to mitigate the Gibbs phenomenon observed in classical Fourier series, Fourier extension has since witnessed extensive research. In recent years, many scholars have made substantial contributions to developing its mathematical foundations, devising efficient algorithms, reduce artifacts, enhance the resolution of inversion results and exploring its applications across various fields Boyd2002A ; Huybrechs2010 ; Lyon2011 ; Matthysen2016FAST ; 2012Sobolev ; Matthysen2018 ; Bruno2022 ; Webb2020 .
Similar to general frame approximations, Fourier extension also encounters challenges due to its near-redundancy. Specifically, it faces the issue of ill-conditioning in the least-squares system used to compute the coefficients of the approximation. Fortunately, accurate Fourier extension approximation can still obtained via suitable regularization. Currently, Truncated Singular Value Decomposition (TSVD) is the commonly used regularization technique, and various fast algorithms have been developed on this basis Lyon2011 ; Matthysen2016FAST ; zhao2024fast . However, solutions obtained by existing methods exhibit oscillations in the extended region, which is detrimental to the application of the extension method in solving partial differential equations (PDEs). A post-processing method described in the literature 2012Sobolev demonstrates that it is possible to find extension functions with desirable properties without compromising accuracy.
According to the extension theorem of Sobolev spaces, it is possible to identify extension functions with desirable properties. This raises a question: Can we find an effective method to directly calculate such extension functions? In fact, when $N$ is finite, $\{e^{\text{i}\frac{k\pi}{T}t}\}_{k=-N}^{N}$ is a basis on [-1,1], that is, the solution to the least squares problem is unique. However, when the TSVD is used, since the singular vectors with singular values $\sigma_{k}\leq\epsilon$ (where $\epsilon$ is the truncation parameter) are discarded, the problem actually becomes a multi-solution. TSVD only yields the solution which minimizes the Euclidean norm of the Fourier coefficients. When a function has a certain smoothness, its Fourier coefficient has a corresponding decay. The smoother the function, the faster the decay rate. Therefore, it is natural to restrict its decay in the selection of solutions. The idea of enforcing special features of regularization approximations has a long history. In his original paper tikhonov1963regularization , Tikhonov (1963) defined the regularization solution $x_{\alpha}$ of the equation
$\mathcal{K}x=y$
as the solution to the following least-squares problem:
$\min\{\|\mathcal{K}x-y\|^{2}+\alpha\|\mathcal{L}x\|^{2}\},$
where $\mathcal{L}$ is a differential operator. Along similar lines, Hansen et. al developed GTSVD and MTSVD methods hansen1989regularization ; hansen1992modified ; hansen1994regularization and Natterer (1984) introduced “regularization in Hilbert scales” natterer1984error . Interestingly, Natterer (1984) has shown that “oversmoothing” (the unknown solution fails to have the expected smoothness) makes no trouble, that is, we can impose stronger constraints than the smoothness of the solution.
In this paper, we will revisit Fourier extension from a new angle: solving compact operator equations with perturbed data. From this perspective, we can more effectively address some concerns previously highlighted in the literature. Furthermore, leveraging results from classical regularization theories, we propose using the weighted best approximate solution as our approximation target. We also employ weighted operators to define the smoothness of the solution, which helps mitigate solution oscillation in the extended region.
1.1 Overview of the paper
In Section 2, we reformulate the Fourier extension, present the theoretical framework of our method and provide a preliminary estimate of the error. In Section 3, we consider the numerical implementation of the algorithm based on uniform sampling, point out the dilemma of the original weighting operator due to machine precision limitations, propose a modified weighting operator, and further test the parameters involved in the algorithm. Section 4 then gives numerical tests of the method to verify the effectiveness of the proposed method. Finally, some conclusions and remarks are presented in Section 5.
2 Reformulation of the Fourier extension
2.1 A new perspective on Fourier extension
Let $T>1$, $\Lambda_{1}=[-1,1]$, $\Lambda_{2}=(-T,T)$, ${L}^{2}(\Lambda_{i})$ and ${H}^{s}(\Lambda_{i})$, $(i=1,2)$ denote the usual Lebesgue and Sobolev spaces, respectively. We consider a system of functions by the restriction of the orthonormal Fourier basis on $\Lambda_{2}$ to $\Lambda_{1}$:
$\Phi=\{\phi_{k}\}_{k\in\mathbb{Z}},\quad\phi_{0}(x)=\frac{1}{\sqrt{2}},\quad\phi_{k}(x)=e^{\text{i}\frac{\pi}{T}kx},k=\pm 1,\pm 2,\ldots,\quad x\in\Lambda_{1}.$
From Huybrechs2010 , this system is a tight frame with frame bound $T$. For $f\in H^{s}(\Lambda_{1})$, we consider the compact operator equation
$\left(\mathcal{F}{{c}}\right)(x):=\sum_{k=-\infty}^{\infty}\hat{c}_{k}\phi_{k}(x)=f(x),\quad x\in\Lambda_{1}.$ (1)
By the extension theorem of the Sobolev space, (1) has multiple solutions and its least square solution and best approximate solution can be defined as follows:
  • ${c}^{\star}\in\ell^{2}$ is called least-squares solution of (1) if
    $\|\mathcal{F}{c}^{\star}-f\|_{L^{2}(\Lambda_{1})}=\inf\{\|\mathcal{F}{c}-f\|_{L^{2}(\Lambda_{1})}~{}|~{}{c}\in\ell^{2}\}$
  • ${c}^{{\dagger}}\in\ell^{2}$ is called the best-approximation solution of of (1) if
[…]
The goal of previous Fourier extension methods can be regarded as obtaining a finite-dimensional approximation of ${c}^{{\dagger}}$.
For any fixed $N$, we define
$\Phi_{N}=\{\phi_{k}\}_{|k|\leq N},$
which is a frame for its span ${H}_{N}=\text{span}(\phi_{N}).$ We define operators
$\mathcal{P}_{N}:\ell^{2}\rightarrow\mathbb{C}^{2N+1},$ ${c}=\{\hat{c}_{k}\}_{k=-\infty}^{\infty}\mapsto{c}_{N}=\{\hat{c}_{k}\}_{|k|\leq N},$
${\mathcal{F}}_{N}:\mathbb{C}^{2N+1}\rightarrow H_{N},$ ${{c}}_{N}=\{\hat{c}_{k}\}_{|k|\leq N}\mapsto\sum_{k=-N}^{N}\hat{c}_{k}\phi_{k},$
${\mathcal{F}}^{*}_{N}:H_{N}\rightarrow\mathbb{C}^{2N+1},$ $f\mapsto\{\langle f,\phi_{k}\rangle\}.$
If we let
${s}_{N}=\mathcal{P}_{N}{c}^{{\dagger}}\quad\hbox{and}\quad f_{N}={\mathcal{F}}_{N}{s}_{N},$
then the Fourier extension method in previous literatures can be formulated as solving the operator equation:
$\mathcal{F}_{N}{{c}}_{N}=f_{N},$ (2)
note that the “exact data” $f_{N}$ is not known, so we can only solve the perturbation equation
$\mathcal{F}_{N}{{c}}_{N}=f,$ (3)
here $f$ is regarded as the “noisy data” of $f_{N}$.
Since $\mathcal{F}_{N}$ is of finite rank, it is a compact operator. Let $(\sigma_{k};v_{k},u_{k})$ be a singular system of $\mathcal{F}_{N}$, then $(\sigma^{2}_{k};v_{k},v_{k})$ be the singular system of $\mathcal{G}=\mathcal{F}^{*}_{N}\mathcal{F}_{N}$. We can use the TSVD method to solve equation (3) and get:
${c}_{N,1}^{\epsilon}=\sum_{\sigma_{k}>\epsilon_{1}}\frac{\langle f,u_{k}\rangle}{\sigma_{k}}v_{k}.$ (4)
Or solve the corresponding normal equation:
$\mathcal{G}{c}=\mathcal{F}_{N}^{*}f$ (5)
and get:
${c}_{N,2}^{\epsilon}=\sum_{\sigma^{2}_{k}>\epsilon_{2}}\frac{\langle\mathcal{F}_{N}^{*}f,v_{k}\rangle}{\sigma^{2}_{k}}v_{k}.$ (6)
Then
$\mathcal{Q}_{N,1}^{\epsilon}f=\mathcal{F}_{N}{c}_{N,1}^{\epsilon}\quad\hbox{or}\quad\mathcal{Q}_{N,2}^{\epsilon}f=\mathcal{F}_{N}{c}_{N,2}^{\epsilon}$ (7)
will be used for the Fourier extension approximation of the function $f$.
From such a perspective, some concerns raised in previous literatures can be easily explained:
  • The oscillation in the extended region is due to the fact that ${c}^{{\dagger}}$ has the smallest norm among all solutions, so it tends to approach zero in the extended region.
  • From the classical regularization theory, we know that the regularization parameter should be related to the error level $\|f-f_{N}\|_{L^{2}(\Lambda_{1})}$, which explains “the coefficient norm first increases exponentially fast, before decaying to $\mathcal{O}(1)$ in the limit.” mentioned in adcock2021frame . When $N$ is small, the error $\|f-f_{N}\|_{L^{2}(\Lambda_{1})}$ is large, and $\epsilon$ is too small for the regularization parameter required for stable solution. As $N$ increases, the error gradually decreases, and after a certain threshold, $\epsilon$ can meet the requirements of stable solution.Therefore, we can actually draw a conclusion: the solutions with large coefficients are unimportant because they mean that $\|f-f_{N}\|_{L^{2}(\Lambda_{1})}$ is large, that is, a desirable approximation to $f$ has not been obtained.
  • From a purely mathematical point of view, there is no difference between (5) and (6). Because as long as $\epsilon_{2}=\epsilon_{1}^{2}$, they are equal. However, in actual calculations, due to the limitation of machine precision, we can only take $\epsilon_{1}=\epsilon_{2}=\epsilon$. In this case, we have
    $\|f-\mathcal{Q}_{N,1}^{\epsilon}f\|\leq\|f-f_{N}\|_{L^{2}(\Lambda_{1})}+\epsilon\|{s}_{N}\|_{\ell^{2}}$
    and
    $\|f-\mathcal{Q}_{N,2}^{\epsilon}f\|\leq\|f-f_{N}\|_{L^{2}(\Lambda_{1})}+\sqrt{\epsilon}\|{s}_{N}\|_{\ell^{2}}.$
    That is to say, if we solve the normal equation, we can only obtain an accuracy of $O(\sqrt{\epsilon})$ instead of $O({\epsilon})$. The literature adcock2019frames also shows that this is true for general frame approximations. Therefore, we should directly focus on the discretization of equation (2) rather than its normal equation.
2.2 Fourier extension towards weighted best-approximation solution
Because $f\in H^{s}(\Lambda_{1})$, we can expect the coefficient $\{\hat{c}_{k}\}$ has a corresponding decay property. Therefore, the following weighted best-approximation solution ${c}^{{\dagger}}_{\mathcal{L}_{p}}$ is a more reasonable goal:
[…]
where
$(\mathcal{L}_{p}{c})_{k}=\rho_{k}\hat{c}_{k}$ (9)
with
$\rho_{k}=:\left\{\begin{aligned} &\sqrt{1+|k|^{2p}},&&0\leq p<\infty,\\ &e^{|k|},&&p=\infty.\end{aligned}\right.$
$p=\infty$ corresponds to the case where $f$ is analytic and it can be seen that when $p=0$, ${c}^{{\dagger}}={c}^{{\dagger}}_{\mathcal{L}_{p}}.$
Remark 2.1
  • Ideally, we would like to take $p=s$, that is, we hope that the solution to the problem has the same smoothness as the function $f$. However, in practice, the value of $s$ may be unknown. In fact, the above operator $\mathcal{L}_{p}$ faces the difficulty of numerical realization due to the limitation of machine precision. We will discuss this in section 3.2 and propose a modified operator.
  • If $p>s$, that is , the smoothness of the solution is lower than the requirement of the operator $\mathcal{L}_{p}$, i. e., no solution of (1) is an element of $\mathcal{D}(\mathcal{L}_{p})$, this is called oversmoothing. The results in natterer1984error ; yu2024super have shown that the regularization method can still obtain the optimal convergence order in the case of oversmoothing (when the regularization parameters are reasonably selected).
Now we set the approximation target the weighted generalized solution ${c}^{{\dagger}}_{\mathcal{L}}$. At this time, the approximation of the solution can be achieved by the “regularization in the Hilbert scale”natterer1984error . In fact, if we let
$\psi_{k}(x)=\frac{1}{\rho_{k}}\phi_{k}(x),\quad\hbox{and}\quad\Psi=\{\psi_{k}\}_{k\in\mathbb{Z}},$
and $\tilde{c}^{{\dagger}}$ be the best approximate solution of the equation
(~$\left(\tilde{\mathcal{F}}\tilde{{c}}\right)(x):=\sum_{k=-\infty}^{\infty}\tilde{c}_{k}\psi_{k}(x)=f(x),\quad x\in\Lambda_{1},$
then we have
${c}^{{\dagger}}_{\mathcal{L}_{p}}=\mathcal{L}_{p}^{-1}\tilde{c}^{{\dagger}}.$
Now we define operator:
$\mathcal{L}_{p,N}:\mathbb{C}^{2N+1}\rightarrow\mathbb{C}^{2N+1},{c}_{N}=\{\hat{c}_{k}\}_{|k|\leq N}\mapsto\tilde{c}_{N}=\{\rho_{k}\hat{c}_{k}\}_{|k|\leq N},$
and let
~$\tilde{\mathcal{F}}_{N}=\mathcal{F}_{N}\mathcal{L}^{-1}_{p,N},\quad{s}_{\mathcal{L},N}=\mathcal{P}_{N}{c}_{\mathcal{L}_{p}}^{{\dagger}}\quad\tilde{s}_{\mathcal{L},N}=\mathcal{L}_{p,N}{s}_{\mathcal{L},N},\quad f_{\mathcal{L},N}=\tilde{\mathcal{F}}_{N}\tilde{s}_{\mathcal{L},N}.$
In this way, obtaining the coefficients of the Fourier extension can be reformulated as solving the equation:
~$\tilde{\mathcal{F}}_{N}\tilde{c}=f_{\mathcal{L},N}$ (10)
Again, the exact right-hand side $f_{\mathcal{L},N}$ is unknown, so we can only solve the perturbation equation:
~$\tilde{\mathcal{F}}_{N}\tilde{c}=f.$ (11)
Obviously, the operator ~$\tilde{\mathcal{F}}_{N}$ is also compact. Assume that its singular system is $(\tilde{\sigma_{k}};{\tilde{v}}_{k},{\tilde{u}}_{k})$ , then the TSVD solution of equation (11) can be given as:
$\tilde{c}_{N}^{\epsilon}=\sum_{\tilde{\sigma}_{k}>\epsilon}\frac{\langle f,\tilde{u}_{k}\rangle}{\tilde{\sigma}_{k}}\tilde{v}_{k}.$ (12)
Now we use
$c_{\mathcal{L},N}^{\epsilon}=\mathcal{L}_{p,N}^{-1}\tilde{c}_{N}^{\epsilon}$ (13)
as the approximation of the coefficient ${s}_{\mathcal{L},N}$ and use
${\mathcal{Q}}_{\mathcal{L},N}^{\epsilon}f=\tilde{\mathcal{F}}_{N}\tilde{c}_{N}^{\epsilon}={\mathcal{F}}_{N}c_{\mathcal{L},N}^{\epsilon}$ (14)
as the Fourier extension approximation of the function $f$.
2.3 Error estimates for $\|f-{\mathcal{Q}}_{\mathcal{L},N}^{\epsilon}f\|$
We now consider the error of the weighted projection ${\mathcal{Q}}_{\mathcal{L},N}^{\epsilon}$.
Theorem 2.1
The weighted TSVD projection ${\mathcal{Q}}_{\mathcal{L},N}^{\epsilon}$ satisfies
$\|f-{\mathcal{Q}}_{\mathcal{L},N}^{\epsilon}f\|_{L^{2}(\Lambda_{1})}\leq\|f-f_{\mathcal{L,N}}\|_{L^{2}(\Lambda_{1})}+\epsilon\|\tilde{s}_{\mathcal{L},N}\|_{\ell^{2}}.$ (15)
Furthermore, if $f\in H^{s}(\Lambda_{1})$, then we have
$\|f-{\mathcal{Q}}_{\mathcal{L},N}^{\epsilon}f\|_{L^{2}(\Lambda_{1})}\leq C\left(N^{-s}+\epsilon\right)\|f\|_{H^{s}(\Lambda_{1})},$ (16)
where $C>0$ is independent of f and N.
Proof
Note that ${\mathcal{Q}}_{\mathcal{L},N}^{\epsilon}$ is the orthogonal projection operator onto $\tilde{H}=\hbox{ span}\{\tilde{u}_{k}:\tilde{\sigma}_{k}>\epsilon\}$, and combined with the triangle inequality, we can get
$\|f-{\mathcal{Q}}_{\mathcal{L},N}^{\epsilon}f\|_{L^{2}(\Lambda_{1})}$$\leq$ $\|f-{\mathcal{Q}}_{\mathcal{L},N}^{\epsilon}f_{\mathcal{L,N}}\|_{L^{2}(\Lambda_{1})}$
$\leq$ $\|f-f_{\mathcal{L,N}}\|_{L^{2}(\Lambda_{1})}+\|f_{\mathcal{L,N}}-{\mathcal{Q}}_{\mathcal{L},N}^{\epsilon}f_{\mathcal{L,N}}\|_{L^{2}(\Lambda_{1})}$
Moreover,
$\|f_{\mathcal{L,N}}-{\mathcal{Q}}_{\mathcal{L},N}^{\epsilon}f_{\mathcal{L,N}}\|_{L^{2}(\Lambda_{1})}$ $=$ $\left\|\sum_{\sigma_{k}\leq\epsilon}\langle f_{\mathcal{L},N},u_{k}\rangle u_{k}\right\|_{L^{2}(\Lambda_{1})}$
$=$ $\left\|\sum_{\sigma_{k}\leq\epsilon}\langle\tilde{\mathcal{F}}_{N}\tilde{\bf s}_{\mathcal{L},N},u_{k}\rangle u_{k}\right\|_{L^{2}(\Lambda_{1})}$
$=$ $\left\|\sum_{\sigma_{k}\leq\epsilon}\sigma_{k}\langle\tilde{\bf s}_{\mathcal{L},N},v_{k}\rangle u_{k}\right\|_{L^{2}(\Lambda_{1})}$
$\leq$ $\epsilon\|\tilde{s}_{\mathcal{L},N}\|$
From the the extension theorem of Sobolev spaces, we know that
$\|\tilde{s}_{\mathcal{L},N}\|_{\ell^{2}}\leq C\|f\|_{H^{s}(\Lambda_{1})}.$
In particular,
$\|f-f_{\mathcal{L,N}}\|_{L^{2}(\Lambda_{1})}\leq\|f-f_{\mathcal{L,N}}\|_{L^{2}(\Lambda_{2})}\leq CN^{-s}\|f\|_{H^{s}(\Lambda_{1})}$
Therefore,
$\|f-{\mathcal{Q}}_{\mathcal{L},N}^{\epsilon}f\|_{L^{2}(\Lambda_{1})}\leq C\left(N^{-s}+\epsilon\right)\|f\|_{H^{s}(\Lambda_{1})}.$
3 Numerical realization based on uniform sampling
3.1 Discretization based on uniform sampling
Selecting different discrete nodes will affect the discrete accuracy of the operator $\mathcal{F}_{N}$, thus having a substantial impact on the calculation results. Previous work distinguishes between uniform sampling, and an optimal sampling set called mapped symmetric Chebychev nodes adcock2014resolution . In this paper, we focus on the case where the sampling points are uniform:
$t_{\ell}=\frac{\ell}{M},\quad\ell=-M,\ldots,M.$
Define the sampling ratio:
$\gamma=M/N,$
after selecting the parameters $N,\gamma,T$, $p$ and let $L={2T\gamma N}$, we get the discrete form of equation (11):
$\tilde{\bf F}_{\gamma,N}^{T,p}\tilde{\bf c}={\bf f},$ (17)
where
$\left(\tilde{\bf F}_{\gamma,N}^{T,p}\right)_{\ell,k}=\frac{1}{\sqrt{L}}\psi_{k}(t_{\ell}),\quad\left({\bf f}\right)_{\ell}=\frac{1}{\sqrt{L}}f(t_{\ell}).$ (18)
Note that if we let
$\left({\bf F}_{\gamma,N}^{T}\right)_{\ell,k}=\frac{1}{\sqrt{L}}\phi_{k}(t_{\ell})$
and ${\bf L}_{p,N}$ be a diagonal matrix with
$({\bf d})_{k}={\rho_{k}},\quad k=-N,\ldots,N.$ (19)
as the diagonal vector, then we have
$\tilde{\bf F}_{\gamma,N}^{T,p}={\bf F}_{\gamma,N}^{T}{\bf L}_{p,N}^{-1}$ (20)
Now let
$N_{\gamma}=\min\{2M+1,2N+1\},$
then the SVD of $\tilde{\bf F}_{\gamma,N}^{T,p}$ can be given as
$\tilde{\bf F}_{\gamma,N}^{T,p}=\sum_{i=1}^{N_{\gamma}}\tilde{\bf u}_{i}\tilde{\bm{\sigma}}_{i}\tilde{\bf v}^{T}_{i},$ (21)
where ${\bm{\sigma}}_{i}$ are the singular values of $\tilde{\bf F}_{\gamma,N}^{T,p}$ such that
$\tilde{\bm{\sigma}}_{1}\geq\ldots\geq\tilde{\bm{\sigma}}_{N_{\gamma}}\geq 0,$
while $\tilde{\bf u}_{i}$ and $\tilde{\bf v}_{i}$ are the left and right singular vectors of $\tilde{\bf F}_{\gamma,N}^{T,p}$, respectively.
Now we can give the TSVD solution of (17) as
$\tilde{\bf c}_{N}^{\epsilon}=\sum_{\tilde{\bm{\sigma}}_{k}>\epsilon}\frac{\langle{\bf f},\tilde{\bf u}_{k}\rangle}{\tilde{\bm{\sigma}}_{k}}{\tilde{\bf v}}_{k}.$ (22)
and get the discrete approximation of $c_{\mathcal{L},N}^{\epsilon}$ and ${\mathcal{Q}}_{\mathcal{L},N}^{\epsilon}f$:
${\bf c}_{{\bf L},N}^{\epsilon}={\bf L}_{p,N}^{-1}\tilde{\bf c}_{N}^{\epsilon},\quad{\bf Q}_{{\bf L},N}^{\epsilon}{\bf f}={\bf F}_{\gamma,N}^{T,p}{\bf c}_{{\bf L},N}^{\epsilon}.$ (23)
3.2 The influence of the parameter $p$ on calculation and the modification of ${\bf L}_{N}$
We first observe the impact of different $p$ values on the singular values of matrix $\tilde{\bf F}_{\gamma,N}^{T,p}$. Fig. 1(a) illustrates the results of the singular values calculated using the SVD functions in Matlab, where $N=200,\gamma=2$ and $T=2$. Similar results are obtained for other combinations of $N$, $\gamma$ and $T$: When $p>0$, some singular values that were originally close to $1$ become decaying, and the decay rate increases with the increase of $p$. When $p$ is too large, the number of singular values greater than the machine precision $\epsilon$ will be significantly reduced. Furthermore, note that the larger $p$ is, the larger the value of $\|\tilde{s}_{\mathcal{L},N}\|$ in the estimate (15) will be, thus affecting the approximation accuracy. In fact, as $p$ increases, we should have chosen a smaller value of $\epsilon$ to achieve approximation, but this is not possible due to the limitation of machine precision.
(a) using $\mathcal{L}_{N}$
(b) using ˘$\breve{\mathcal{L}}_{N}$ with $K_{0}=20$
Figure 1: The singular values of $\tilde{\bf F}_{\gamma,N}^{T,p}$ for various $p$.
(a) $T=4$, $p=4$
(b) $\omega=20$, $T=2$
(c) $\omega=20$, $p=4$
Figure 2: The approximation error against $N$ for various $\omega$, $T$, and $p$ using ${\mathcal{L}}_{N}$.
Next, we take
$f(t)=\exp(\text{i}\pi\omega t)$
to conduct a preliminary test on the performance of the algorithm. Here and what follows, the approximation error evaluated as the maximum pointwise error over an equi-sampled grid that is ten times denser than that used for construction. Unless otherwise specified, the value of $\epsilon$ is $141e-$ 14. Fig. 2 shows how the error changes with $N$ when the parameters $\omega$, $T$, and $p$ take different values. It is evident that these three parameters significantly affect the calculation accuracy, with their increase leading to a reduction in accuracy. This is because $\exp(\text{i}\pi\omega x)$ corresponds to $\phi_{k}$ when $k=\omega T$, so we can obtain the estimate of $\|\tilde{s}_{\mathcal{L},N}\|$ in (15) as:
$\|\tilde{s}_{\mathcal{L},N}\|_{\ell^{2}}\sim\left\{\begin{aligned} &\left(\omega T\right)^{p},&0\leq p<\infty,\\ &e^{\omega T},&p=\infty.\end{aligned}\right.$ (24)
This means that if the approximated function contains high-frequency terms that cannot be ignored, the approximation error will be large, which is not what we expect. Therefore, we consider making the following corrections to the operator $\mathcal{L}_{N}$: Define the modification operator ˘$\breve{\mathcal{L}}_{N}$ as
˘$\breve{\mathcal{L}}_{N}:\mathbb{C}^{2N+1}\rightarrow\mathbb{C}^{2N+1},{c}_{N}=\{\hat{c}_{k}\}_{|k|\leq N}\mapsto\tilde{c}_{N},$ (25)
when $0\leq p<\infty$
$\left(\tilde{c}_{N}\right)_{k}=\left\{\begin{aligned} &\frac{1}{K_{0}}\hat{c}_{k},&&|k|<K_{0},\\ &\left[{\left(\frac{|k|-K_{0}}{(N-K_{0})^{\frac{K_{0}}{K_{0}+C}}T}\right)^{p}}+\frac{1}{K_{0}}\right]\hat{c}_{k},&&|k|\geq K_{0},\end{aligned}\right.$ (26)
and when $p=\infty$
$\left(\tilde{c}_{N}\right)_{k}=\left\{\begin{aligned} &\frac{1}{K_{0}}\hat{c}_{k},&&|k|<K_{0},\\ &\left[\exp\left(\frac{|k|-K_{0}}{(N-K_{0})^{\frac{K_{0}}{K_{0}+C}}T}\right)-\frac{K_{0}}{K_{0}+1}\right]\hat{c}_{k},&&|k|\geq K_{0},\end{aligned}\right.$ (27)
where $K_{0}$ and $C$ are constants to be determined. Such corrections are reasonable because the goal of introducing the weighted generalized inverse is essentially to characterize the attenuation of the coefficients, which is an asymptotic property when $k$ tends to infinity. The parameter $K_{0}$ describes the starting point of the function attenuation, which can be estimated during the calculation process, as follows: First, replace ${\bf L}_{N}$ with $\breve{{\bf L}}_{N}(p=4,K_{0}=0)$ in (22)–(23) to calculate ${\bf c}_{\breve{\bf{L}},N}^{\epsilon}$ and then select $K_{0}$ as
$K_{0}=\inf\left\{K~{}|\left|\left({\bf c}_{\breve{\bf L},N}^{\epsilon}\right)_{k}\right|>\varepsilon,\forall\,\,|k|>K\right\}.$ (28)
In which, $\varepsilon$ is a specified threshold. For calculations using the correction operator ˘$\breve{\mathcal{L}}_{N}$, we only need to modify the diagonal vectors in (19) accordingly and continue with (21)–(23). The correction plays the role of reducing the norm of $\breve{\bf L}_{N}{\bf c}_{\breve{\bf L},N}^{\epsilon}$ while maintaining the coefficient ${\bf c}_{\breve{\bf L},N}^{\epsilon}$ decay. In subsequent experiments, we always take $C=32$ and $\varepsilon=1e-1$. We also tested other values, and the calculation results are not sensitive to their values. Fig 1(b) shows the singular values of the matrix $\tilde{\bf F}_{\gamma,N}^{T,p}$ when the correction operator ˘$\breve{\mathcal{L}}_{N}$ is used. It can be seen that the rate of decrease with $p$ is controlled, so that we can obtain more effective singular vectors for calculation within the range allowed by the machine precision. Fig. 3 shows how the error changes with $N$ when the parameters $\omega$, $T$, and $p$ take different values after the above correction. It can be seen that the defects have been effectively corrected.
(a) $T=4$, $p=4$
(b) $\omega=20$, $T=2$
(c) $\omega=20$, $p=4$
Figure 3: The approximation error against $N$ for various $\omega$, $T$, and $p$ using ˘$\breve{\mathcal{L}}_{N}$.
(a) $\gamma=2$, $p=2$, $N=200$
(b) $\omega=20$, $p=2$, $N=200$
(c) $\omega=20$, $\gamma=2$, $N=200$
(d) $\omega=20$, $p=2$,$\gamma=2$
Figure 4: The approximation error $\|f-\tilde{{\bf Q}}_{\breve{\bf L},N}^{\epsilon}{\bf f}\|$ against $T$ for different values of $\omega$, $\gamma$ and $p$.
3.3 The influence of the parameter $T$ on calculation accuracy
We still use $f(t)=\exp(\text{i}\pi\omega t)$ to test the algorithm. Figure 4 shows the results of the approximation error against the parameter $T$ when other parameters have different values. Three distinct regions are visible in all graphs in Fig. 4:
  • Region $I_{1}:=\{T:1<T<{T}_{1}\}$, where the error decreases rapidly to the machine accuracy as $T$ increases.
  • Region $I_{2}:=\{T:{T}_{1}<T<{T}_{2}\}$, where the error is maintained close to the machine accuracy.
  • Region $I_{3}:=\{T>{T}_{2}\}$, where the error rebounds, and the increase in $T$ will cause the error to increase instead.
It can be seen that the value of $T_{1}$ is influenced by $\gamma$, but not by $\omega$, $N$ and $p$, while $T_{2}$ is influenced by $\gamma$, $N$ and $\omega$. In fact, when $N$ is fixed, increasing $\gamma$ can improve the discretization accuracy of $\tilde{\bf F}_{\gamma,N}^{T,p}$ to ~$\tilde{\mathcal{F}}_{N}$, which can also be achieved by increasing the value of $T$. This explains that the approximation accuracy is affected by the interaction of the two parameters. We tested the approximate values of f $T_{1}$ for different $\gamma$ and listed them in Table 1. We can see that the value of the product $T_{1}\gamma$ is very close. For $T_{2}$, note that $\exp(\text{i}\pi\omega t)$ corresponds to the $k=\omega T$ term in $\{\phi_{k}\}$, so $f$ can only be effectively approximated when $N\geq T\omega$, that is,
$T_{2}=\frac{N}{\omega}.$ (29)
In actual calculations, (29) tells us that when approximating functions of the same frequency, an increase in $T$ means that a larger $N$ is required, and a larger $\gamma$ corresponds to a larger $M$, which will affect the final calculation efficiency. This is not the focus of this paper. For a fast algorithm based on boundary intervals data, please refer to zhao2024fast .
Table 1: The approximation value of ${T}_{1}$ for various $\gamma$.
$\gamma=1$ $\gamma=2$ $\gamma=3$ $\gamma=4$ $\gamma=8$
$T_{1}$ $5.6$ $2.3$ $1.6$ $1.2$ $1.08$
4 Numerical tests
In this section, some numerical experiments are given to verify the proposed method. All the tests were computed on a Windows 10 system with a 16 GB memory, Intel(R) Core(TM)i7-8500U CPU1.80GHz using MATLAB 2016b. Through the analysis in the previous section, the basic calculation process of the final algorithm can be given as follows:
  • Input parameters $N$, $\gamma$, $p$, $T$ and data $f(t_{\ell})$;
  • Estimate $K_{0}$ according to (28);
  • Replace ${\bf L}_{N}$ with $\breve{{\bf L}}_{N}$ in (22)–(23) to calculate ${\bf c}_{\breve{\mathcal{L}},N}^{\epsilon}$ and ${\bf Q}_{\breve{\bf L},N}^{\epsilon}{\bf f}$.
The method proposed in this paper is mainly to overcome the oscillation of the solution in the extended region, so we pay more attention to the parameter $p$. The subsequent numerical experiments mainly show the impact of different values of $p$ on the solution, and the parameters $\gamma$ and $T$ are uniformly set to $\gamma=3$ and $T=2$.
We show the approximation results of the following six test functions:
$f_{1}(t)=t^{2},\quad f_{2}(t)=\exp(t),\quad f_{3}(t)=e^{\sin(2.7\pi t)+\cos(\pi t)},$
$f_{4}(t)=\text{erf}(10t),\quad f_{5}(t)=\cos(10t^{2}),\quad f_{6}(t)=\cos\left(\frac{10}{1+25t^{2}}\right).$
Fig. 5 shows the weighted projection ${{\bf Q}}_{\breve{\bf L},N}^{\epsilon}{\bf f}$ of $f_{i}$, $i=1,2,\ldots 6$. From the previous analysis, we know that $p=0$ corresponds to the original method, so it shows the characteristic of oscillation around 0 in the extension region. When $p$ takes other values, the function has different forms in the extension region, which is related to the properties of the approximated function. It can be seen that the proposed method effectively alleviates the high-frequency oscillation of the original method in the extension region.
(a) $f_{1}(t)$
(b) $f_{2}(t)$
(c) $f_{3}(t)$
(d) $f_{4}(t)$
(e) $f_{5}(t)$
(f) $f_{6}(t)$
Figure 5: Calculation results of ${{\bf Q}}_{\breve{\bf L},N}^{\epsilon}{\bf f}$ for the various test functions ($N=200$).
Figs. 6-8 show how the approximation errors of different test functions and their first-order and second-order derivatives vary with $N$. The derivatives are calculated by trigonometric interpolation using the FFT on $[-T,T]$. It can be seen that the approximation errors of the functions are very close for different values of $p$, and the proposed method has a significant improvement on the errors of the first-order and second-order derivatives. We also noticed that after $N$ reaches a certain threshold, the error of the derivative calculation tends to get worse as $N$ increases, which is caused by the method we use to calculate the derivatives.
(a) $f_{1}(t)$
(b) $f_{2}(t)$
(c) $f_{3}(t)$.
(d) $f_{4}(t)$
(e) $f_{5}(t)$
(f) $f_{6}(t)$
Figure 6: Approximation errors for the various test functions.
(a) $f_{1}(t)$
(b) $f_{2}(t)$
(c) $f_{3}(t)$
(d) $f_{4}(t)$
(e) $f_{5}(t)$
(f) $f_{6}(t)$
Figure 7: Approximation error of the first-order derivative for various test functions.
(a) $f_{1}(t)$
(b) $f_{2}(t)$
(c) $f_{3}(t)$
(d) $f_{4}(t)$
(e) $f_{5}(t)$
(f) $f_{6}(t)$
Figure 8: Approximation error of the second-order derivative for various test functions.
5 Conclusions and remarks
We conclude the paper with some final remarks.
  • We present a new approach for calculating Fourier extensions, which can effectively mitigate the issue of oscillations in the extended region. This idea can also be applied to other types of frame approximations.
  • This paper does not address computational efficiency. For those interested in a fast algorithm, reference [21] provides details on the implementation process when $p=0$. The value of $p$ does not have a substantial effect on the algorithm itself.
  • The design of the weighted operator $\breve{\mathcal{L}}$ is an issue worthy of further consideration, which we will explore in future work.
  • Existing Fourier extensions consider the overall approximation of the function, and the singularity of the function at any position can substantially impact the approximation result. The results of this paper and zhao2024fast actually point out the potential possibility of piecewise approximation, which will be the focus of our further work.
Data Availability
Derived data supporting the findings of this study are available from the authors on request.